Katie is willing to pay $3 for a pound of apples. She buys 2 pounds of apples for $2 per pound. The lowest price the seller would accept is $1.50 per pound. How much value is created as a result of this trade?
Answer #1

Katie's willingness to pay (WTP) is $3 but Katie buys at p=$2. So the consumer surplus is (3-2)*2=$2 for 2 pounds of apple.

Seller's willingness to sell (WTS) is $1.5 but he gets $2. So the producer surplus is (2-1.5)*2=$1 for 2 pounds of apple.

Therefore total value created inthis trade is (2+1)=$3